OpenStudy (anonymous):

Evaluate tanu-cot3u/u-pi/4 as u approaches pi/4

OpenStudy (anonymous):

I have to use l hopitals rule here.

OpenStudy (anonymous):

I know that its in the form 0/0 and its in ratio form so i can take the derivative. \[\lim_{u \rightarrow \pi/4} \frac{ tanu-\cot3u }{ u-\pi/4 } = \frac{f \prime}{g \prime} = \sec^2u+3\csc^23u\]

OpenStudy (anonymous):

from here my issue is figuring out how to evaluate the function. How is wolfram alpha getting 2 for sec^2 at u = pi/4

OpenStudy (anonymous):

and 2 for 3csc^23u?